OpenStudy (anonymous):

find the limit of : limx->pi/2 3sinx/2cosx+4

myininaya (myininaya):

\[\frac{3 \sin(\frac{\pi}{2})}{2\cos(\frac{\pi}{2})+4}=\frac{2(1)}{2(0)+4}=\frac{2}{4}=\frac{1}{2}\]

OpenStudy (anonymous):